# Towards Visual Text Grounding of Multimodal Large Language Model

### v1

@@ Abstract
Despite the existing evolution of Multimodal Large Language Models (MLLMs), a non-neglectable limitation remains in their struggle with visual text grounding, especially in text-rich images of documents. Document images, such as scanned forms and infographics, highlight critical challenges due to their complex layouts and textual content. However, current benchmarks do not fully address these challenges, as they mostly focus on visual grounding on natural images, rather than text-rich document images. Thus, to bridge this gap, we introduce TRIG, a novel task with a newly designed instruction dataset for benchmarking and improving the Text-Rich Image Grounding capabilities of MLLMs in document question-answering. Specifically, we propose an OCR-LLM-human interaction pipeline to create $800$ manually annotated question-answer pairs as a benchmark and a large-scale training set of $90$k synthetic data based on four diverse datasets. A comprehensive evaluation of various MLLMs on our proposed benchmark exposes substantial limitations in their grounding capability on text-rich images. In addition, we propose two simple and effective TRIG methods based on general instruction tuning and plug-and-play efficient embedding, respectively. By finetuning MLLMs on our synthetic dataset, they promisingly improve spatial reasoning and grounding capabilities.
@@ 1 Introduction
Despite the remarkable advancements in LLMs [64, 65, 4, 53] and MLLMs [85, 16, 41], the trustworthiness of their generated outputs remains a critical concern [40, 60, 20]. While these models can produce fluent and coherent responses, they often lack grounding capability, which can lead to potential hallucinations [21, 58, 70, 35]. Grounding capability is defined as the model’s ability to accurately localize relevant regions in the visual content based on the provided semantic description [51, 45, 83, 22, 82]. This capability is essential for ensuring that the responses are accurate and verifiable.
@@
Existing grounding efforts in MLLMs mainly focus on natural images, where the task involves associating textual descriptions with corresponding visual elements, such as objects or scenes[57, 7, 87, 89, 56, 82]. However, there is a significant gap in the literature when it comes to visual text grounding in text-rich document images. Document images, such as scanned forms, charts, and complex posters, present unique challenges that differ markedly from those found in natural images [50, 48, 49, 86]. They often feature a mix of textual and graphical elements and require precise localization and understanding of content, especially textual content. An example is shown in Figure 1, for the given text-rich document image, we expect LLMs to not only generate the answer alone but also to provide the corresponding grounded bounding boxes that can support its answer, which requires deeper spatial understanding and reasoning capabilities, and sometimes better instruction-following abilities to correctly provide the grounding information in the desired formats.
@@
Figure 1: An example from InfographicsVQA. Question: What is the resource and expertise level needed for Pinterest? Answer: high. The LLM is expected to generate the answer together with the corresponding grounded bounding boxes that can support its answer, which requires deeper spatial understanding and reasoning and sometimes instruction-following abilities.
@@
Despite its importance, there is no established task specifically designed to evaluate the visual text grounding capabilities of MLLMs on Text-Rich Document Question-Answering tasks and no well-formulated corresponding instruction datasets and baseline methods for the purpose. The absence of such a well-formulated task limits the ability to systematically assess and improve the performance of different models in this domain. Thus to bridge this gap, we introduce TRIG, a novel Text-Rich Image Grounding task with instruction set for document QA grounding, along with its corresponding benchmark notated as TRIG-Bench. TRIG-Bench consists of $800$ question-answer pairs manually collected from DocVQA [50], ChartQA [48], InfographicsVQA [49], and TRINS datasets [86], along with human-inspected ground-truth bounding boxes that support the answer to the corresponding question. The TRIG instruction dataset contains approximately $90$k training instances generated and verified by the GPT4o model from the above-mentioned data sources. They provide a standardized framework for evaluating the ability of MLLMs to accurately ground their responses to document-related visual questions.
@@
For text-rich document images, we mainly focus on visual texts as the main grounding target. Considering the supreme performance of modern OCR models [8, 59, 13, 32, 31, 73, 61] and the promising reasoning ability of current LLMs like GPT4o, an OCR-LLM-human interaction pipeline is proposed for the benchmark construction. Specifically, for every given VQA pair, we first utilize PaddleOCR to detect and recognize all the texts. Given all these OCR results, LLMs are asked to judge which bounding boxes support the answer to the corresponding question, followed by another LLM evaluating the correctness of the chosen grounded bounding boxes. The resulting data has already been of good quality, as it is generated and verified by the powerful GPT4o model. Next, more human participants will manually inspect the generated results and add the correct ones to build our benchmark data.
@@
In addition, two types of methods are further proposed, including the general instruction-tuning-based method and a novel embedding-based method, as potential baselines for TRIG. Specifically, for the instruction-tuning-based method, following the implementation of modern MLLM instruction tuning [38], we train an MLLM with supreme document grounding capability. Moreover, we further propose an efficient plug-and-play embedding-based method, whose inference efficiency is highly promising due to the unnecessity of the iterative token generation process, which largely speeds up the grounding process though with lower performance scores. Our main contributions:
@@
  • We introduce a novel task text-rich document image grounding (TRIG), its corresponding benchmark TRIG-Bench and an instruction tuning dataset. The dataset and benchmark are the first of their kind and provide a standardized framework for evaluating MLLMs in this domain, filling a critical gap in existing research.
@@ 2 Related Works
Grounding [18] is essential for effective human communication with machines. Currently, grounding datasets have been utilized for vision-language pre-training and improvement for both object-level recognition [26] and language acquisition [47]. Recent studies propose to integrate text and grounding regions into token sequences [76, 44, 66, 69, 68] within language modeling frameworks. Building on such approaches, researchers have developed a series of grounded MLLMs, including GPT4ROI [87], Kosmos-2 [55], Shikra [7], PVIT [5], BuboGPT [90], Qwen-VL [1], and Ferret [81]. Despite their impressive performance, these models mainly focus on the grounding of natural image objects, and the grounding of text-rich document images is still under-explored. For the two most related works that also mentioned text grounding on document images, P2G [6] includes the OCR model into the QA pipeline as an information amplifier; TG-Doc [71] utilizes grounding capability to improve models’ QA capabilities. Both of these works treat the text grounding process as an intermediate step and do not conduct any evaluation of the grounding capability. On the contrary, we entirely focus on the grounding capability and first propose the benchmark for it. Another related work is TextMonkey [42], which describes a task that is similar to ours, but there are no examples, illustrations, and systematic evaluations of this ability included. This situation further showcases the value of our benchmark.
@@ 4 Proposed Methods
Along with the benchmark and corresponding training data, we propose two baseline methods for this task: (1) the instruction-tuning-based [38, 88, 33] method, (we use the term instruction-based for simplicity), that treats this grounding task as an instruction-following task by iterative next token prediction, and (2) the efficient embedding-based method that directly finds the image patches that have greatest similarities with the input text embeddings.
The instruction-based method formulates the generation of bounding boxes into a token-prediction framework, in which the inputs include a curated prompt $x$ and an image, and the output will be a sequence of tokens as the response $y$. For simplicity, we denote each data sample as a triplet $(\text{Image},x,y)$ and let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In this type of method, the bounding box coordinates are also represented by sets of tokens. For simplicity, we denote the text string that represents each bounding box as $b_{i}$ for the $i$th bounding box. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by minimizing the next-token-prediction loss on each data $(\text{Image},x,y)$.
@@ 4.1 Instruction-Based Method
For the OCR-free Grounding setting, i.e., evaluation setting 1, the input prompt contains only the original question and the instruction guiding LLMs to generate formatted grounded bounding boxes, which can be simply noted as $x$. The output of this setting includes the answer $y$ to the question concatenated with the texts of corresponding supportive bounding boxes, which can be denoted as $y^{\prime}$. Thus, the loss function in this setting can be denoted as:
@@
where $G$ is the number of golden grounded bounding boxes, and [;][;][ ; ] represents the concatenation of strings.
For the OCR-based Grounding setting, i.e. the evaluation setting 2, all the bounding boxes generated by the OCR module will be provided in the input prompt, thus the new input $x^{\prime}$ in this setting will be the concatenation of the original question, the instruction guiding LLMs to select the supportive bounding boxes, and all the bounding boxes. The output of this setting is identical to the previous setting, which includes the answer $y$ to the question concatenated with the texts of corresponding supportive bounding boxes, which can be denoted as $y^{\prime}$. Thus, the loss function in this setting can be denoted as:
$L_{\theta}=-\sum_{j=1}^{l}\log p_{\theta}\left(y^{\prime}_{j}|\text{Image},x^{\prime},y^{\prime}_{<j}\right),x^{\prime}=[x;b_{1};..;b_{N}],y^{\prime}=[y;b^{\prime}_{1};..;b^{\prime}_{G}]$ (2)
@@ 4.2 Embedding-Based Method
In addition to the above instruction-based method, motivated by ColPali [14], an embedding-based method is also provided. For instruction-based methods, the instruction-following capability of the LLMs severely affects the final performance, as shown in our experimental results, most of the existing LLMs are not capable of following complex user-customized instructions like ours, leading to unsatisfactory performances. However, the embedding-based method disentangles the grounding capability from the instruction-following capability and makes this process much more efficient than continuous next-token prediction. Thus, it can serve as a plug-and-play module adaptive to MLLMs without grounding capability. An illustrative figure for our embedding-based method pipeline is shown in Figure 4. Different from the instruction-based method requiring MLLMs to generate tokens iteratively, which is extremely time-consuming for a large number of forward processes, the embedding-based method only needs two times of forward processes, which largely increases the efficiency of the grounding process.
@@
During the inference phase, the main goal is to locate the image patches that have the highest similarities with the text embeddings, aligning with the training objective, the text is the concatenation of the question and the ground-truth answer. Unlike tokens that are naturally discrete, the image patches are generated by cutting the continuous image, thus, the information of one object might be unintentionally contained by several adjacent patches, further causing the low similarities of each patch. To alleviate this phenomenon, a novel adjacent patch embedding merging technique is first implemented, in which each image patch embedding $H^{\text{Image}}_{i}$ will be updated as the average of it and its surrounding embeddings:
@@
The above similarity vector $S$ represents the similarity of each image patch to the whole text embeddings. We assume the image patches with higher similarities contain the information mentioned in the text, and thus can be used as the grounding area. Then, the image patches can be found by selecting the top-$k$ highest values. Moreover, to further improve the robustness and alleviate the severe influence of the selection of $k$, a further 2-level selection framework is utilized. In the framework, we first select the image patches with top-$5$ similarities, and then progressively include later top-$k$ patches if it is in the surroundings of the top-$5$ patches.
After obtaining the image patches, for the OCR-free Grounding setting, the performance is directly calculated by utilizing the pixels in the selected image patches, and for the OCR-based Grounding setting, the bounding boxes that overlap with image patches are selected as the final bounding boxes for calculating the performance.
@@ 5 Experimental Results
For the instruction-based method, we utilize LLaVA-v1.5-Vicuna-13B [36] as our base model to be fine-tuned. For the embedding-based method, we utilize PaliGemma-3B [3] as our base model to be fine-tuned. Detailed training configurations, data statistics, and examples can be found in the supplementary material. All experiments are performed on Nvidia A100 GPUs.
@@ 5.1 Main Results
The overall evaluation results across various MLLMs are presented in Table 1 (Evaluation Setting 1), Table 2 (Evaluation Setting 2). The MLLMs we evaluated are listed as follows: LLaVA-v1.6-Vicuna-13B, LLaVA-v1.6-Vicuna-7B [36, 38, 37], Phi3-V[62], DeepSeek-VL-7B-chat [43], Idefics2-8B [24, 25], Qwen-VL [2], CogVLM2-Llama3-19B [67], InternLM-XComposer2-VL-7B [10], InternLM-XComposer2-4KHD-7B [11], Monkey-Chat [34], MiniCPM-Llama3-V 2.5 [77], and GPT series.
As shown in Table 1, all the existing models, including open-source models and the powerful GPT4o, do not perform well for the OCR-free Grounding setting, in which they are required to generate supportive bounding boxes without any additional bounding box information. Under this circumstance, both our proposed methods outperform GPT4o by a large margin: our embedding-based method reaches the average IoU of $10.0$, and our instruction-based method reaches the average IoU of $29.98$, compared with $1.51$ for GPT4o. Several conclusions can be made from the evaluation setting 1 results: (1) The capability of existing models to generate grounded bounding boxes from scratch is limited, and thus, the training specifically for this setting is required. (2) Our proposed method is effective, and the quality of our corresponding training dataset is promising as our performance outperforms the powerful GPT4o by a large margin. (3) Instruction tuning on reasonably good LLMs can yield much better performance than utilizing the embedding-based method, while the embedding-based method is much more efficient during inference: Specifically, the instruction method takes an average of $2.37$ seconds per document, whereas the Embedding Method takes $0.74$ seconds per document, calculated based on inference on $1$ Nvidia A100 GPU.
@@
Qwen-VL 4.36 4.33 28.00 7.27 1.07 1.06 8.25 1.83 0.70 0.71 4.50 1.20 1.51 1.83 3.32 2.08 4.50
@@
As shown in Table 2, for the OCR-based Grounding setting, most of the models can reach a much better performance as the OCR information has been provided. However, the performances of open-source models still have a large gap with GPT4o. Both of our methods achieve much better performances compared with the existing open-source models while failing to beat GPT4o, which is due to its inherently powerful instruction-following capability for customized instructions.

### v2

@@ Abstract
Despite the existing evolution of Multimodal Large Language Models (MLLMs), a non-neglectable limitation remains in their struggle with visual text grounding, especially in text-rich images of documents. Document images, such as scanned forms and infographics, highlight critical challenges due to their complex layouts and textual content. However, current benchmarks do not fully address these challenges, as they mostly focus on visual grounding on natural images, rather than text-rich document images. Thus, to bridge this gap, we introduce TRIG, a novel task with a newly designed instruction dataset for benchmarking and improving the Text-Rich Image Grounding capabilities of MLLMs in document question-answering. Specifically, we propose an OCR-LLM-human interaction pipeline to create $800$ manually annotated question-answer pairs as a benchmark and a large-scale training set of 9090k synthetic data based on four diverse datasets. A comprehensive evaluation of various MLLMs on our proposed benchmark exposes substantial limitations in their grounding capability on text-rich images. In addition, we propose two simple and effective TRIG methods based on general instruction tuning and plug-and-play efficient embedding, respectively. By finetuning MLLMs on our synthetic dataset, they promisingly improve spatial reasoning and grounding capabilities.
@@ 1 Introduction
Despite the remarkable advancements in LLMs and MLLMs, the trustworthiness of their generated outputs remains a critical concern [41, 59]. While these models can produce fluent and coherent responses, they often lack grounding capability, which can lead to potential hallucinations [22, 58, 36]. Grounding capability is defined as the model’s ability to accurately localize relevant regions in the visual content based on the provided semantic description [51, 45, 77, 23, 76]. This capability is essential for ensuring that the responses are accurate and verifiable.
@@
Existing grounding efforts in MLLMs mainly focus on natural images, where the task involves associating textual descriptions with corresponding visual elements, such as objects or scenes[57, 9, 80, 82, 56, 76, 33, 35]. However, there is a significant gap in the literature when it comes to visual text grounding in text-rich document images. Document images, such as scanned forms, charts, and complex posters, present unique challenges that differ markedly from those found in natural images [50, 48, 49, 79, 7, 6]. They often feature a mix of textual and graphical elements and require precise localization and understanding of content, especially textual content. An example is shown in Figure 1, for the given text-rich document image, we expect LLMs to not only generate the answer alone but also to provide the corresponding grounded bounding boxes that can support its answer, which requires deeper spatial understanding and reasoning capabilities, and sometimes better instruction-following abilities to correctly provide the grounding information in the desired formats.
@@
Figure 1: An example from InfograhicsVQA. Question: What is the resource and expertise level needed for Pinterest? Answer: high. The LLM is expected to generate the answer together with the corresponding grounded bounding boxes that can support its answer, which requires deeper spatial understanding and reasoning, and sometimes instruction-following abilities.
@@
Despite its importance, there is no established task specifically designed to evaluate the visual text grounding capabilities of MLLMs on Text-Rich Document Question-Answering tasks and no well-formulated corresponding instruction datasets and baseline methods for the purpose. The absence of such a well-formulated task limits the ability to systematically assess and improve the performance of different models in this domain. Thus to bridge this gap, we introduce TRIG, a novel Text-Rich Image Grounding task with instruction set for document QA grounding, along with its corresponding benchmark notated as TRIG-Bench. TRIG-Bench consists of $800$ question-answer pairs manually collected from DocVQA [50], ChartQA [48], InfographicsVQA [49], and TRINS datasets [79], along with human-inspected ground-truth bounding boxes that support the answer to the corresponding question. The TRIG instruction dataset contains approximately 9090k training instances generated and verified by the GPT4o model from the above-mentioned data sources. They provide a standardized framework for evaluating the ability of MLLMs to accurately ground their responses to document-related visual questions.
@@
For text-rich document images, we mainly focus on the visual texts on them as the main grounding target. Considering the supreme performance of modern OCR models and the promising reasoning ability of current LLMs like GPT4o, an OCR-LLM-Human interaction pipeline is proposed for the benchmark construction. Specifically, for every given VQA pair, we first utilize PaddleOCR to detect and recognize all the texts. Given all these OCR results, LLMs are asked to judge which bounding boxes support the answer to the corresponding question, followed by another LLM evaluating the correctness of the chosen grounded bounding boxes. The resulting data has already been of good quality, as it is generated and verified by the powerful GPT4o model. Next, more human participants will manually inspect the generated results and add the correct ones to build our benchmark data.
@@
In addition, two types of methods are further proposed, including the general instruction-tuning-based method and a novel embedding-based method, as potential baselines for TRIG. Specifically, for the instruction-tuning-based method, following the implementation of modern MLLM instruction tuning [39], we train an MLLM with supreme document grounding capability. Moreover, we further propose an efficient plug-and-play embedding-based method, whose inference efficiency is highly promising due to the unnecessity of the iterative token generation process, which largely speeds up the grounding process, though with lower performance scores. Our main contributions:
@@
  • We introduce a novel task, text-rich document image grounding (TRIG), its corresponding benchmark TRIG-Bench, and an instruction tuning dataset. The dataset and benchmark are the first of their kind and provide a standardized framework for evaluating MLLMs in this domain, filling a critical gap in existing research.
@@ 2 Related Works
Grounding [20] is essential for effective human communication with machines. Currently, grounding datasets have been utilized for vision-language pre-training and improvement for both object-level recognition [28] and language acquisition [47]. Recent studies propose to integrate text and grounding regions into token sequences [70, 44, 64] within language modeling frameworks. Building on such approaches, researchers have developed a series of grounded MLLMs, including GPT4ROI [80], Kosmos-2 [55], Shikra [9], PVIT [5], BuboGPT [83], Qwen-VL [1], and Ferret [75]. Despite their impressive performance, these models mainly focus on the grounding of natural image objects, and the grounding of text-rich document images is still underexplored. For the two most related works that also mentioned text grounding on document images, P2G [8] includes the OCR model into the QA pipeline as an information amplifier; TG-Doc [66] utilizes grounding capability to improve models’ QA capabilities. Both of these works treat the text grounding process as an intermediate step and do not conduct any evaluation of the grounding capability. On the contrary, we entirely focus on the grounding capability and first propose the benchmark for it. Another related work is TextMonkey [42], which describes a task that is similar to ours, but there are no examples, illustrations, or systematic evaluations of this ability included. This situation further showcases the value of our benchmark.
@@ 4 Proposed Methods
Along with the benchmark and corresponding training data, we propose two baseline methods for this task: (1) the instruction-tuning-based [39, 81] method (we use the term instruction-based for simplicity), that treats this grounding task as an instruction-following task by iterative next token prediction, and (2) the efficient embedding-based method that directly finds the image patches that have the greatest similarities with the input text embeddings.
The instruction-based method formulates the generation of bounding boxes into a token-prediction framework, in which the inputs include a curated prompt $x$ and an image, and the output will be a sequence of tokens as the response $y$. For simplicity, we denote each data sample as a triplet $(\text{Image},x,y)$ and let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In this type of method, the bounding box coordinates are also represented by sets of tokens. For simplicity, we denote the text string that represents each bounding box as $b_{i}$ for the $i$th bounding box. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by minimizing the next-token-prediction loss on each data $(\text{Image},x,y)$:
@@ 4.1 Instruction-Based Method
For the OCR-free Grounding setting, i.e., the evaluation setting 1, the input prompt contains only the original question and the instruction guiding LLMs to generate formatted grounded bounding boxes, which can be simply noted as $x$. The output of this setting includes the answer $y$ to the question concatenated with the texts of corresponding supportive bounding boxes, which can be denoted as $y^{\prime}$. Thus, the loss function in this setting can be denoted as:
@@
where $G$ is the number of golden grounded bounding boxes, and $[;]$ represents the concatenation of strings.
For the OCR-based Grounding setting, i.e., the evaluation setting 2, all the bounding boxes generated by the OCR module will be provided in the input prompt, thus, the new input $x^{\prime}$ in this setting will be the concatenation of the original question, the instruction guiding LLMs to select the supportive bounding boxes, and all the bounding boxes. The output of this setting is identical to the previous setting, which includes the answer $y$ to the question concatenated with the texts of corresponding supportive bounding boxes, which can be denoted as $y^{\prime}$. Thus, the loss function in this setting can be denoted as:
$L_{\theta}$ $=-\sum_{j=1}^{l}\log p_{\theta}\!\left(y^{\prime}_{j}\mid\text{Image},\,x^{\prime},\,y^{\prime}_{<j}\right),$ (2)
$x^{\prime}$ $=[x;b_{1};\dots;b_{N}],\qquad y^{\prime}=[y;b^{\prime}_{1};\dots;b^{\prime}_{G}].$
@@ 4.2 Embedding-Based Method
In addition to the above instruction-based method, motivated by ColPali [17], an embedding-based method is also provided. For instruction-based methods, the instruction-following capability of the LLMs severely affects the final performance, as shown in our experimental results. Most of the existing LLMs are not capable of following complex user-customized instructions like ours, leading to unsatisfactory performances. However, the embedding-based method disentangles the grounding capability from the instruction-following capability and makes this process much more efficient than continuous next-token prediction. Thus, it can serve as a plug-and-play module that is adaptive to MLLMs without grounding capability.
@@
During the inference phase, the main goal is to locate the image patches that have the highest similarities with the text embeddings, aligning with the training objective, the text is the concatenation of the question and the ground-truth answer. Unlike tokens that are naturally discrete, the image patches are generated by cutting the continuous image; thus, the information of one object might be unintentionally contained in several adjacent patches, further causing the low similarities of each patch. To alleviate this phenomenon, a novel adjacent patch embedding merging technique is first implemented, in which each image patch embedding $H^{\text{Image}}_{i}$ will be updated as the average of it and its surrounding embeddings:
@@
The similarity vector $S$ represents the similarity of each image patch to the whole text embedding. The image patches with higher similarities contain the information mentioned in the text and thus can be used as the grounding area. Then, the image patches can be found by selecting the top-$k$ highest values. Moreover, to further improve the robustness and alleviate the severe influence of the selection of $k$, a further 2-level selection framework is utilized. In the framework, we first select the image patches with top-$5$ similarities and then progressively include later top-$k$ patches if it is in the surroundings of the top-$5$ patches. The detailed ablation studies can be found in the supplementary material.
@@ 5 Experimental Results
Qwen2-VL-2B 0.00 0.00 0.00 0.00 0.00
Qwen2-VL-7B 0.00 0.00 0.00 0.00 0.00
Qwen2.5-VL-3B 0.00 0.00 0.00 0.00 0.00
Qwen2.5-VL-7B 0.00 0.00 0.00 0.00 0.00
LLaVA-OV-1B 0.00 0.00 0.00 0.00 0.00
LLaVA-OV-7B 0.00 0.00 0.00 0.00 0.00
InternVL2-1B 0.00 0.00 0.00 0.00 0.00
InternVL2-8B 0.00 0.00 0.00 0.00 0.00
InternVL2.5-1B 0.00 0.00 0.00 0.00 0.00
InternVL2.5-8B 0.00 0.00 0.00 0.00 0.00
@@
For the instruction-based method, we utilize LLaVA-v1.5-Vicuna-13B [37] as our base model to be fine-tuned. For the embedding-based method, we utilize PaliGemma-3B [4] as our base model to be fine-tuned. Detailed training configurations can be found in the supplementary material.
@@ 5.1 Main Results
The overall evaluation results across various MLLMs are presented in Table 1 (Evaluation Setting 1), Table 2 (Evaluation Setting 2). The MLLMs we evaluated are listed as follows: LLaVA-v1.6-Vicuna-13B, LLaVA-v1.6-Vicuna-7B [37, 39, 38], Phi3-V[60], DeepSeek-VL-7B-chat [43], Idefics2-8B [25, 26], Qwen-VL [2], CogVLM2-Llama3-19B [65], InternLM-XComposer2-VL-7B [14], InternLM-XComposer2-4KHD-7B [15], Monkey-Chat [34], MiniCPM-Llama3-V 2.5 [71], InternVL2 [10], LLaVA-OV [27], InternVL2.5 [10], Qwen2.5-VL [3], Qwen2-VL [63], Gemini-2-flash [12], Gemini-2.5-flash [13], and GPT series.
As shown in Table 1, all the existing models, including open-source models and the powerful GPT4o, not perform well for the OCR-free Grounding setting, in which they are required to generate supportive bounding boxes without any additional bounding box information. Under this circumstance, both our proposed methods outperform GPT4o by a large margin: our embedding-based method reaches the average IoU of $10.0$, and our instruction-based method reaches the average IoU of $29.98$, compared with $1.51$ for GPT4o. Several conclusions can be made from the evaluation setting 1 results: (1) The capability of existing models to generate grounded bounding boxes from scratch is limited, and thus, the training specifically for this setting is required. (2) Our proposed method is effective and the quality of our corresponding training dataset is promising as our performance outperforms the powerful GPT4o by a large margin. (3) Instruction tuning on reasonably good LLMs can yield much better performance than utilizing the embedding-based method, while the embedding-based method is much more efficient during inference: Specifically, the instruction method takes an average of $2.37$ seconds per document, whereas the Embedding Method takes $0.74$ seconds per document, calculated based on inference on $1$ Nvidia A100 GPU.
@@
InternVL2-8B 0.07 0.50 0.07 0.13 3.92 4.62 6.33 4.50 1.11 1.63 1.42 1.36 5.60 6.39 6.39 6.28 3.10
Qwen-VL-7B 4.36 4.33 28.00 7.27 1.07 1.06 8.25 1.83 0.70 0.71 4.50 1.20 1.51 1.83 3.32 2.08 4.50
Qwen2-VL-2B 5.21 6.65 11.56 7.86 3.48 3.67 10.25 5.25 0.49 0.75 1.00 0.84 7.23 8.13 11.17 8.79 4.52
LLaVA-OV-1B 3.57 4.45 7.90 5.36 4.24 4.13 9.88 5.59 0.51 0.75 1.08 0.86 7.05 8.21 10.79 8.65 4.86
Qwen2.5-VL-3B 3.77 5.74 5.37 5.02 2.46 3.00 3.58 3.02 0.10 0.17 0.17 0.17 10.19 12.75 14.59 12.34 5.05
InternVL2.5-VL-1B 3.01 2.91 16.19 4.70 0.56 0.55 5.25 0.99 0.03 0.03 0.50 0.06 14.47 16.12 33.54 19.20 8.16
@@
As shown in Table 2, for the OCR-based Grounding setting, most of the models can reach a much better performance with the OCR information provided. However, the performances of open-source models still have a large gap with GPT4o. Both of our methods achieve better performances compared with the existing open-source models while failing to beat GPT4o, which is due to its inherently powerful instruction-following capability for customized instructions.